OpenStudy (anonymous):

A particular type of bacteria grows at a rate of 35 % per day. If my dog got infected with 9900 bacterium when her leg was cut by barbed wire, how many bacterium will be present in 11 days when I finally decide I need to take the dog to the veterinarian.

OpenStudy (cwrw238):

this is a progression which increases by a factor of 1.35 everyday after 1 day the number of bacteria = 9900 * 1.35 after 2 days the number of bacteria = 9900 * 1.35 ^2 after 3 .. .. .............................. = 9900 *1.35^3 and so on do you see the pattern? so what do you think the expression for the number of bacteria will be after 11 days?

OpenStudy (anonymous):

9900 * 1.35^11 The answer is extremely long. How do I input into my webwork?

OpenStudy (cwrw238):

use your calculator

OpenStudy (cwrw238):

or the calculator on your computer

OpenStudy (cwrw238):

your answer should be 268,724

OpenStudy (cwrw238):

- poor dog!

OpenStudy (anonymous):

Thank you! I was pressing the wrong button on my calculator. I had the right formula, but you clarified it for me. Yes, and poor dog.

OpenStudy (cwrw238):

yw